What is the wavelength and frequency of light required to ionize a hydrogen atom (that is, to completely remove its electron)? This light is in what part of the EM spectrum?
Answer by Alan3354(69443)

What is the wavelength and frequency of light required to ionize a hydrogen atom (that is, to completely remove its electron)? This light is in what part of the EM spectrum?
=================================
The ionization energy of hydrogen is 1312 kJ/mol. The energy required to ionize a single hydrogen atom is this quantity divided by Avagadro's number (the number of particles in a mole):
(1.312*10^6 J/mol)/(6.022*10^23 atoms/mol) =
2.179*10^-18 J/atom

The energy of a photon is related to its wavelength by:
E = h*c/wavelength
where c is the speed of light in a vacuum = 2.998*10^8 m/s
h is Planck's constant = 6.626*10^-34 (kg*m^2)/s
Solving for the wavelength of a photon that has enough energy to ionize hydrogen:
wavelength = (2.998*10^8 m/s)*(6.626*10^-34 (kg*m^2)/s)/(2.179*10^-18 J)
wavelength = 9.118*10^-8 m = 91.18 nm